OpenStudy (anonymous):

Which of the following is a solution to the equation log6 x + log6 (x – 1) = 1 ? Which of the following is a solution to the equation log6 x + log6 (x – 1) = 1 ? @Mathematics

OpenStudy (anonymous):

x = 4 x = 6 / 5 x = 2 x = 3

OpenStudy (anonymous):

\[\log_{a}x +\log_{a} y = \log_{a} x.y \] thats why it will be \[\log_{6} x.(x-1) =1\] x.(x-1) =6^1 x^2 - x =6 x^2-x-6=0 (x-3).(x+2)=0 x=3 or x=-2 so answer is x=3